Jolexa:

A professor determined the relationship between the time spent studying (in hours) and performance on an exam. Performance = 70.443 + 4.885 * (time) Ann studied 2.6 hours for the last exam. However there was a concert in town the night before and her score was 16 points lower than expected. What was her score on this exam, rounded to the nearest integer?

dude:

You want to know the actual grade, you are given the equation for (expected) performance First, substitute the 2.5 hours into time \(P = 70.443+4.885\times\color{green}{2.6}\) From then, you subtract what you got from above and 16 (because Ann got 16 points lower) \(P-16=?\)